Grade 12General Physics

Electric dipole moment of CuSo4 molecule is 3.2×10^-32 Cm.find the separation between copper and sulphate ions?

Mohit Jaat Gatiyan
9 Years agoGrade 12

1 Answer

Khimraj
9 Years ago
Dipole moment P = q*l where Q is magnitude of charge on one particle of dipole which is 2e in case of Cu+2 so distance between them is given by l= P/2e where e is charge on electron = 1.6 *10^-19C so distance is 10^-13 m